The sum of two numbers is 26. Find the two numbers if the larger number is 2 less than three times the smaller number.
Answer by kchande(5)

let us say 2 numbers are x and y and y is the larger number.
Since their sum is 26, so x + y = 26
also y = 3x-2 based on the 2nd fact.
Now from Ist equation y = 26 -x
equate the 2 y,
26 -x = 3x -2
so 4x = 28
x = 7
now y = 26 - 7 = 19
so numbers are 7 and 9.
Thanks,
kchande